# ASME Letters in Dynamic Systems and Control

### v1

@@ Abstract
In this paper, we address the stabilization problem for force‐controlled nonholonomic mobile robots under safety‐critical constraints. We propose a continuous, time‐invariant control law based on the $\gamma m$-quadratic programming ($\gamma m$-QP) framework, which unifies control Lyapunov functions (CLFs) and control barrier functions (CBFs) to enforce both stability and safety in the closed‐loop system. For the first time, we construct a global, time-invariant, strict Lyapunov function for the closed‐loop nonholonomic mobile robot system with a nominal stabilization controller in polar coordinates; this strict Lyapunov function then serves as the CLF in the QP design. Next, by exploiting the inherent cascaded structure of the vehicle dynamics, we develop a CBF for the mobile robot via an integrator backstepping procedure. Our main results guarantee both asymptotic stability and safety for the closed-loop system. Both the simulation and experimental results are presented to illustrate the effectiveness and performance of our approach.
@@ 1 Introduction
In [18], a safety-critical controller is designed for connected automated vehicles, where the vehicles are modeled by integrators. Using the double‐integrator model, the multi‐agent collision avoidance problem has been studied via CBF approaches in [6]. Based on the first‐order unicycle (kinematic) model, CBF‐based obstacle avoidance has been addressed in [19, 20]. In particular, in [19], a CBF backstepping approach is proposed for the kinematic unicycle model. In [20], an obstacle‐avoidance strategy for nonholonomic‐integrator vehicles is proposed by regulating vehicle speed and orientation separately via two CBFs while maintaining nonzero forward speed in dynamic environments using velocity obstacles. However, none of these existing works provides guarantees of asymptotic stability for the closed-loop system. Moreover, a more realistic model for vehicle applications is to consider the second-order full dynamical (kinematics-kinetics) model of the unicycle [21, 22]. However, to the best of the authors’ knowledge, few studies have addressed the stabilization problem for force-controlled nonholonomic vehicles subject to safety-critical constraints.
In this paper, we address the stabilization problem for force-controlled nonholonomic mobile robots under safety-critical constraints. We propose a continuous, time‐invariant control law based on the $\gamma m$-QP framework to enforce both stability and safety in the closed‐loop system.
The main contributions of this work include:
  1. (i)
    For the first time, we construct a global, time-invariant, strict Lyapunov function for the closed‐loop nonholonomic mobile robot system with a nominal stabilization controller in polar coordinates. This strict Lyapunov function then serves as the CLF in the $\gamma m$-QP design.
  2. (ii)
    We present experimental results to validate the effectiveness of the proposed approach and to demonstrate the performance of the developed controller. The experiments show that the proposed method is applicable to scenarios such as autonomous parking with obstacle avoidance and inter-vehicle collision avoidance.
@@
In our previous work [4], we construct a strict Lyapunov function for the closed-loop mobile robot with a nominal stabilization controller in the large (i.e., on any compact subset of the state space), where it serves as a CLF in the safety-critical control design. However, the constructed Lyapunov function is not global, meaning that it depends on the initial configuration of the vehicle. To the best of the authors’ knowledge, a global, time-invariant, strict Lyapunov function has not yet been reported in the literature. Moreover, the problem of eliminating potential undesired equilibria, e.g., via introducing additional constraints in the QP [23], remains out of scope of this Letter.
@@ 2 Preliminaries on Safety-Critical Control
where the state $x\in\mathbb{R}^{n}$ and the control $u\in\mathbb{R}^{m}$. We assume that $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ are locally Lipschitz and $f(0)=0$. Recall that a $C^{\infty}$ function $V:\mathbb{R}^{n}\to\mathbb{R}_{\geq 0}$ is said to be a (global) CLF for (1), if $V$ is positive definite, proper, and satisfies the following implication:
@@ 4.1 Simulation Results
All parameters are given in SI units. The initial conditions of the robot are randomly selected as $(x_{0},y_{0},\theta_{0})=(-3.15,2.96,-1.43)$, and the robot is initially at rest. To illustrate the effectiveness of the proposed approach, three controllers (i.e., nominal controller (25), CLF-QP (12), CLF-CBF-QP (19)) were implemented and compared. We assume that a circular obstacle is located at (1,0)(-1,0) with radius $r=0.3$. That is, the admissible set is given by $C_{0}:=\{(x,y)\in\mathbb{R}^{2}:h_{0}(x,y)=40((x+2)^{2}+y^{2}-0.3^{2})\}$. We define $\alpha:=\frac{\mu W\dot{W}|_{f_{\text{nom}}}}{2(W+1)^{2}}-\frac{1}{2}|\zeta|^{2}$ and $\alpha_{h}(s):=2s$. The control parameters are set to $\lambda=3$, $k_{\rho}=2$, $k_{\alpha}=2$, $k_{z}=4$, $k_{\omega}=4$, $\mu=0.05$, $l_{v}=1$, $l_{\omega}=1$ and $m=1$. The simulation results are shown in Figs. 1-2, which demonstrate that the proposed CLF-CBF $\gamma m$-QP controller effectively achieves parking with obstacle avoidance.
@@ Acknowledgment
The work of was supported in part by GSoE at CCNY, and in part by the PSC-CUNY Award, jointly funded by The Professional Staff Congress and The City University of New York.

### v2

@@ Abstract
In this paper, we address the stabilization problem for force-controlled nonholonomic mobile robots under safety-critical constraints. We propose a continuous, time-invariant control law based on the $\gamma m$-quadratic programming ($\gamma m$-QP) framework, which unifies control Lyapunov functions (CLFs) and control barrier functions (CBFs) to enforce both stability and safety in the closed‐loop system. For the first time, we construct a global, time-invariant, strict Lyapunov function for the closed-loop nonholonomic mobile robot full-dynamic system with a nominal stabilization controller in polar coordinates; this strict Lyapunov function then serves as the CLF in the QP design. Next, by exploiting the inherent cascaded structure of the vehicle dynamics, we develop a CBF for the mobile robot via an integrator backstepping procedure. Our main results guarantee both asymptotic stability and safety for the closed-loop system. Both the simulation and experimental results are presented to illustrate the effectiveness and performance of our approach.
@@ 1 Introduction
Constructions of global strict Lyapunov functions for nonholonomic systems are by no means trivial or straightforward. To the best of the authors’ knowledge, only a few studies have presented such global and strict Lyapunov functions for nonholonomic mobile robots [3]. Very recently, global strict Lyapunov functions were developed for the kinematic unicycle model using backstepping and forwarding techniques in [5, 6].
@@
In [20], a safety-critical controller is designed for connected automated vehicles, where the vehicles are modeled by integrators. Using the double‐integrator model, the multi‐agent collision avoidance problem has been studied via CBF approaches in [8]. Based on the first‐order unicycle (kinematic) model, CBF‐based obstacle avoidance has been addressed in [21, 22]. In particular, in [21], a CBF backstepping approach is proposed for the kinematic unicycle model. In [22], an obstacle‐avoidance strategy for nonholonomic‐integrator vehicles is proposed by regulating vehicle speed and orientation separately via two CBFs while maintaining a nonzero forward speed in dynamic environments using velocity obstacles. However, none of these existing works provides guarantees of asymptotic stability for the closed-loop system. Moreover, a more realistic model for vehicle applications is to consider the second-order full dynamical (kinematics-kinetics) model of the unicycle [23, 24]. However, to the best of the authors’ knowledge, few studies have addressed the stabilization problem for force-controlled nonholonomic vehicles subject to safety-critical constraints.
In this paper, we address the stabilization problem for force-controlled nonholonomic mobile robots under safety-critical constraints. We propose a continuous, time-invariant control law based on the $\gamma m$-QP framework to enforce both stability and safety in the closed‐loop system. The main contribution of this work is that, for the first time, we construct a global, time-invariant, and strict Lyapunov function for the closed-loop nonholonomic mobile robot (kinematic–dynamic) system under a nominal stabilization controller in polar coordinates. This global strict Lyapunov function then serves as the CLF in the $\gamma m$-QP design. Compared with our previous result in [4], where the Lyapunov function is valid only within an arbitrarily large but still bounded region, the present work develops a globally valid Lyapunov function whose definition and derivative properties hold for all admissible states, without any restriction on the region of attraction. Compared with the designs in [5, 6], our controller is much simpler and applied to the full dynamic unicycle system. We also present both simulation and experimental results to demonstrate the performance of the developed controller. In particular, the experimental results show that the proposed method is applicable to scenarios such as autonomous parking with obstacle avoidance and inter-vehicle collision avoidance.
@@
In our previous work [4], we construct a strict Lyapunov function for the closed-loop mobile robot with a nominal stabilization controller in the large (i.e., on any compact subset of the state space), where it serves as a CLF in the safety-critical control design. However, the constructed Lyapunov function is not global, meaning that it depends on the initial configuration of the vehicle. To the best of the authors’ knowledge, a global, time-invariant, strict Lyapunov function has not yet been reported in the literature for the full kinematic-kinetic mobile robot system. Moreover, the problem of eliminating potential undesired equilibria, e.g., via introducing additional constraints in the QP [25], and explicit robustness analyses—such as the treatment of saturation effects or the derivation of input-to-state stability (ISS) bounds, remain out of scope of this Letter.
@@ 2 Preliminaries on Safety-Critical Control
where the state $x\in\mathbb{R}^{n}$ and the control $u\in\mathbb{R}^{m}$. We assume that $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ are locally Lipschitz functions and $f(0)=0$. Recall that a $C^{\infty}$ function $V:\mathbb{R}^{n}\to\mathbb{R}_{\geq 0}$ is said to be a (global) CLF for (1), if $V$ is positive definite, proper, and satisfies the following implication:
@@ 3.1 Construction of the global CLF
Remark 1.
It is necessary to point out the difference between the global Lyapunov function (12) and the Lyapunov function we developed in our previous work [4]. In [4], the Lyapunov function is valid within an arbitrarily large, but still bounded region. The formulation in [4] could guarantee safety-critical stabilization, but the parameters of the associated Lyapunov function need to be chosen according to the initial conditions. In contrast, the Lyapunov function developed in the present paper is globally valid, which eliminates the need for any parameter adjustment based on initial states, thereby providing a rigorous analytical foundation for safety-critical control across the full admissible space.
Remark 2.
The new velocity coordinate $z:=(v-v^{*})/{\rho}$ is used in the dynamic error equation (25). However, one must verify the boundedness of the term $\dot{v}^{*}(t)$ in (25) along trajectories because this term is used in the control law (26). Direct calculations show that
$\frac{{\rm d}v^{*}}{{\rm d}t}$ $=$ $k_{\rho}(\cos(\alpha)\dot{\rho}-\sin(\alpha)\dot{\alpha}\rho)$
$=$ $k_{\rho}\left(-\cos^{2}(\alpha)v-\sin(\alpha)\rho\left(\frac{v}{\rho}\sin{\alpha}-\omega\right)\right).$
It can be observed that the term $1/\rho$ cancels out, and thus the boundedness of $\dot{v}^{*}(t)$ follows.
Remark 3.
[…]
@@ 3.2 Construction of the CBF
Remark 4.
It can be seen from (16) and (9) that the constructed CBF $h$ has a relative degree of $\{1,1\}$ for all states satisfying $(v,\omega)\neq(0,0)$. On the set $(v,\omega)=(0,0)$}, it holds that $L_{F}h=0$ and $\alpha_{h}(h)\geq 0$; hence, the safety property still holds. This means that even in those states where $u$ has no influence on $\dot{h}$, the system remains safe. This is intuitively clear because, in such cases, the mobile robot simply stops (i.e., $(v,\omega)=(0,0)$). With the CBF $h$, the robot may safely stop in front of an obstacle while maintaining the safety property. This behavior is reminiscent of the undesirable equilibria observed in CLF–CBF–based QPs [25] and may be solved by introducing additional QP constraints [25]. However, the problem of eliminating potential undesired equilibria remains out of the scope of this Letter.
@@ 4.1 Simulation Results
All parameters are given in SI units. The initial conditions of the robot are randomly selected as $(x_{0},y_{0},\theta_{0})=(-3.15,2.96,-1.43)$, and the robot is initially at rest. The target position is set at the origin. To illustrate the effectiveness of the proposed approach, three controllers (i.e., nominal controller (26), CLF-QP (12), CLF-CBF-QP (20)) were implemented and compared. We assume that a circular obstacle is located at (2,0)(-2,0) with radius $r=0.3$. That is, the admissible set is given by $C_{0}:=\{(x,y)\in\mathbb{R}^{2}:h_{0}(x,y)=40((x+2)^{2}+y^{2}-0.3^{2})\}$. We define $\alpha:=\frac{\mu W\dot{W}|_{f_{\text{nom}}}}{2(W+1)^{2}}-\frac{1}{2}|\zeta|^{2}$ and $\alpha_{h}(s):=2s$. The control parameters are set to $\lambda=3$, $k_{\rho}=2$, $k_{\alpha}=2$, $k_{z}=4$, $k_{\omega}=4$, $\mu=0.05$, $l_{v}=1$, $l_{\omega}=1$ and $m=1$. The simulation results are shown in Figs. 1-2, which demonstrate that the proposed CLF-CBF $\gamma m$-QP controller effectively achieves parking with obstacle avoidance.
@@ 4.2 Experimental Results
Remark 5.
In the experiments, a safety margin was deliberately introduced to account for the physical size of the mobile robot. Specifically, the effective obstacle boundaries used in the CBF ($r=0.2$) were defined to be slightly larger than the actual physical obstacles. This ensures that even when the robot’s center of mass reaches the boundary of the defined obstacle, the robot’s physical body still remains clear of contact with the real obstacle.
@@ Acknowledgment
The first author would like to express his gratitude to Prof. Miroslav Krstić of UCSD for the fruitful discussions regarding the construction of the global strict Lyapunov function and Remarks 2 and 3. The authors would also like to thank the Associate Editor and the anonymous reviewers for their useful suggestions.
The work of Bo Wang was supported in part by GSoE at CCNY, and in part by the PSC-CUNY Award, jointly funded by The Professional Staff Congress and The City University of New York.